OpenStudy (anonymous):

keep getting stuck on this problems -s+4=7s-3

OpenStudy (anonymous):

Ok so first your main goal is to get s by itself

OpenStudy (anonymous):

First- I would add s to both sides -s+4=7s-3 4=8s-3

OpenStudy (anonymous):

Then, add three to both sides

OpenStudy (anonymous):

7=8s

OpenStudy (anonymous):

now divide each side by 8 s=7/8

OpenStudy (anonymous):

Does this make sense?

OpenStudy (anonymous):

alittle

OpenStudy (anonymous):

What don't you understand? Maybe I can explain it a little better.

OpenStudy (anonymous):

oka

OpenStudy (anonymous):

like I said your ultimate goal is to get s by itself. first you have to get you s's on one side of the = sign and your other numbers on the other side So I said.... -s+4=7s-3 what is the easiest way to combine the -s and the 7s?

OpenStudy (anonymous):

ooo i get it

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

you add s to both sides then add the 3 to both sides 7=8s 7/8=s